# Towards User-Focused Research in Training Data Attribution for Human-Centered Explainable AI

### v1

@@ Acknowledgements.
- Liao et al. (2020) Q. Vera Liao, Daniel Gruen, and Sarah Miller. 2020. Questioning the AI: Informing Design Practices for Explainable AI User Experiences. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–15. https://doi.org/10.1145/3313831.3376590
@@
- Nauta et al. (2023) Meike Nauta, Jan Trienes, Shreyasi Pathak, Elisa Nguyen, Michelle Peters, Yasmin Schmitt, Jörg Schlötterer, Maurice van Keulen, and Christin Seifert. 2023. From Anecdotal Evidence to Quantitative Evaluation Methods: A Systematic Review on Evaluating Explainable AI. ACM Comput. Surv. 55, 13s, Article 295 (jul 2023), 42 pages. https://doi.org/10.1145/3583558
@@
- Patnaik and Becker (1999) Dev Patnaik and Robert Becker. 1999. Needfinding: The Why and How of Uncovering People’s Needs. Design Management Journal (Former Series) 10, 2 (1999), 37–43. https://doi.org/10.1111/j.1948-7169.1999.tb00250.x arXiv:https://onlinelibrary.wiley.com/doi/pdf/10.1111/j.1948-7169.1999.tb00250.x
@@
- Pezeshkpour et al. (2022) Pouya Pezeshkpour, Sarthak Jain, Sameer Singh, and Byron Wallace. 2022. Combining Feature and Instance Attribution to Detect Artifacts. In Findings of the Association for Computational Linguistics: ACL 2022. Association for Computational Linguistics, Dublin, Ireland, 1934–1946. https://doi.org/10.18653/v1/2022.findings-acl.153
@@
- Rong et al. (2023) Yao Rong, Tobias Leemann, Thai-Trang Nguyen, Lisa Fiedler, Peizhu Qian, Vaibhav Unhelkar, Tina Seidel, Gjergji Kasneci, and Enkelejda Kasneci. 2023. Towards Human-Centered Explainable AI: A Survey of User Studies for Model Explanations. IEEE Trans. Pattern Anal. Mach. Intell. 46, 4 (nov 2023), 2104–2122. https://doi.org/10.1109/TPAMI.2023.3331846
@@
- Schioppa et al. (2022) Andrea Schioppa, Polina Zablotskaia, David Vilar, and Artem Sokolov. 2022. Scaling Up Influence Functions. Proceedings of the AAAI Conference on Artificial Intelligence 36, 8 (Jun. 2022), 8179–8186. https://doi.org/10.1609/aaai.v36i8.20791
- Shen et al. (2023) Hua Shen, Chieh-Yang Huang, Tongshuang Wu, and Ting-Hao Kenneth Huang. 2023. ConvXAI: Delivering Heterogeneous AI Explanations via Conversations to Support Human-AI Scientific Writing. In Companion Publication of the 2023 Conference on Computer Supported Cooperative Work and Social Computing (Minneapolis, MN, USA) (CSCW ’23 Companion). Association for Computing Machinery, New York, NY, USA, 384–387. https://doi.org/10.1145/3584931.3607492
@@
- Turri et al. (2024) V. Turri, K. Morrison, K. Robinson, C. Abidi, J. Forlizzi, A. Perer, and R. Dzombak. 2024. Transparency in the Wild: Navigating Transparency in a Deployed AI System to Broaden Need-Finding Approaches. ACM Conference on Fairness, Accountability, and Transparency (ACM FAccT) (2024).
- von Thienen et al. (2019) Julia P. A. von Thienen, William J. Clancey, and Christoph Meinel. 2019. Theoretical Foundations of Design Thinking. Springer International Publishing, Cham, 13–38. https://doi.org/10.1007/978-3-319-97082-0_2
@@
- Wang et al. (2024a) Fulton Wang, Julius Adebayo, Sarah Tan, Diego Garcia-Olano, and Narine Kokhlikyan. 2024a. Error discovery by clustering influence embeddings. In Proceedings of the 37th International Conference on Neural Information Processing Systems (New Orleans, LA, USA) (NIPS ’23). Curran Associates Inc., Red Hook, NY, USA, Article 1809, 13 pages.
- Wang et al. (2024b) Sheng-Yu Wang, Aaron Hertzmann, Alexei A Efros, Jun-Yan Zhu, and Richard Zhang. 2024b. Data Attribution for Text-to-Image Models by Unlearning Synthesized Images. arXiv preprint arXiv:2406.09408 (2024).
@@ Appendix A Inclusion criteria: High-risk application areas
Appendix B Interview study themes
The thematic coding analysis of the survey study (N=10) yielded 29 themes corresponding to six theme areas and four unexpected themes (§ 4.2). In the following, we further elaborate on these themes.
B.1. Theme area: Role of ML systems
The theme area “Role of ML systems” groups three themes that represent what kind of role ML systems take in our participant’s work:
Work assistant: ML systems act as work assistants as they take workload off the participant. Chatbots generally fill the role of a work assistant that “[takes work off of the participant’s hands and makes their work easier]” (P10) and is “available around the clock” (P10). Participants use chatbot systems to improve their writing in English (P1, P11), to search for information where previously they would ”[ask] Google” (P1), or to ideate research ideas (P11). Moreover, P10 trusts their company-internal chatbot enough to redirect simple employee questions to the chatbot.
Decision support: As decision support systems, ML systems deliver information that acts as a basis for decisions taken by end-users, e.g. diagnosis support (P6). This is the main role taken by the ML systems that the developer participants work on (“[We] make predictions about where we should do our [pharmaceutical] studies and our development.” (P8), “try to accommodate the policymakers so that they’re able to decide [where more health services are needed]” (P9)). P3 uses ML systems to identify and explain the contributing factors to product issues: “If we can predict it, we can also have an idea what are the factors mostly creating this phenomenon.”
Automation: In some cases, ML systems are deployed for the automation of some tasks, e.g. autonomous driving (P2) or automated processing of camera recordings (P4).
B.2. Theme area: Workflow with ML systems
The theme area “Workflow with ML systems” groups four themes that emerged in the interviews related to standard or often-occuring workflows:
Collaboration with domain experts: Developers and system users collaborate closely for building and evaluating ML models, where bugs are reported to the developers by the end-users (P2, P3, P6, P4, P9). This collaboration is especially strong if the ML system is developed in-house, and less strong if the ML system is a product the company purchased.
Standard model choices: Developers rarely build their own model architectures but opt for standard ones (“[If] it’s already an established problem, you’re probably not going to do better than an algorithm that’s already been laid out to solve that problem for you.” (P8)).
Data-focused debugging process: The development work in practice is centred around data. Developer participants mention that most often, data artefacts are the reason for bugs in the model (P2, P3, P4, P7, P8, P9).
Data quality checks: Data quality checks are a set part of the data preprocessing pipeline (P2, P3, P4, P8). P2 and P9 reported that they first assess data quality before inspecting the model in debugging. This includes looking for any possible data artefacts, especially missing data.
B.3. Theme area: Role of training data
The theme area “Role of training data” groups five themes stemming from participant’s responses when talking about the role of training data in the interviews:
Most important variable in a model: Model developer participants consistently view training data as the most important variable in a model, e.g. “[we] […] believe that […] the models can only be as good as the […] data that you feed in.” (P4). Hence, the debugging process is focused around data (Theme: Data-focused debugging process).
Interesting but no capacity: This theme applies to system user participants, in particular P6. They mention that they would be curious about training data to understand whether the model learned from high-quality data but “[it] is a luxury that [requires time]”, highlighting the practical constraint of time pressure.
Not interesting: This theme applies to system user participants. They were unaware of which training data may have been used to train the models (P1, P10) and mentioned that this does not matter too much for their work as they cannot change it anyhow.
Increasing training data as a bug fix: A reoccuring theme in our interviews is that training data is flexible – the training set can be expanded. This is often done in practice when encountering a bug, e.g. P2 explained that collecting more data is a common way to overcome model shortcomings in autonomous driving.
Training data artefacts as main bug causes: In line with previous themes highlighting the importance of training data to the development process, participants explicitly mention that most often, training data artefacts are the cause of bad model performance. Hence, data artefacts are also the first thing that developers look our for when encountering a bug.
B.4. Theme area: Challenges in working with ML systems
The theme area “Challenges in working with ML systems” groups six themes related to the main pain points that participants mentioned
Trust calibration: Our findings agree with Kim et al. (2023): It is unclear how much and when a system can be trusted. P1 sometimes finds themselves in a dilemma in which they wish to learn something from the chatbot, but are unable to calibrate their trust in the response due to missing knowledge: “I don’t know everything regarding this topic. I [don’t even] know what he’s replying to me.” (P1).
Validation and Evaluation: P2 mentions that the validation itself is a challenge due to multiple requirements that the ML system should fulfil. In practice, the predictive performance of a model matters, along with other objectives such as inference speed, whether the model is bias-free, and more. In addition, A data-centric aspect of evaluation being challenging is the difficulty of evaluation when labels are absent.
Data quality issues: Data quality issues are a major challenge when working with ML systems. In Table 2, we detail the quality issues mentioned by the pariticipants.
Resource constraints: For developer participants, memory and compute constraints are relevant in the development and evaluation process of their models, especially to P2 and P4 as they work with ML systems on the edge.
Stochastic nature of ML models: The stochastic nature of ML models poses a specific challenge to deep models in the development process and debugging tasks: “[The] same data set, same model, you train multiple times, you can get [different] results.” (P2). It is therefore difficult to understand if a bug was truly fixed or if it is a consequence of randomness.
Misuse due to lacking know-how: P10 mentions inadequate know-how in ML system usage as a challenge: “the employees often don’t manage to ask the chatbot the right questions”.
B.5. Theme area: Use of (feature attribution) XAI in practice
The theme area “Use of (feature attribution) XAI in practice” groups six themes related to how XAI is already being used by the participants.
XAI is not used: Some system user participants were unaware of XAI since explanations are not a part of the ML tools they use (P1, P10, P11).
XAI is little used: This theme is related to the system users. P6 reports that XAI tools they used in radiology images so far (i.e. heatmaps) do not deliver a full answer to the why question, as counterfactual information is missing: “[If] I just get an overall highlight in these basal lung regions and the prediction that is atelectasis, I still don’t know why this is atelectasis but not pleural effusion or consolidation.” Hence, they only use it when the information is useful.
Model development: For developer participants, XAI is most commonly used as a tool for model development (P2, P3, P8). As such, XAI tools (within our participants, these are exclusively feature attribution explanations, i.e., SHAP (Lundberg and Lee, 2017) or visualisation of attention) offer explanations for per-example debugging of e.g. wrong predictions (“we work with hands-on examples that we then […] relate the input features that take a poor or a good prediction [to].” - P8).
Understand real-world phenomena: P3 described the use of XAI as a tool to understand phenomena represented by the ML model: “[Building] the model, the whole purpose is to get some explainability. Because […] we know that [a problem is] happening and predicting doesn’t really add value. But if we can predict it, we can also have an idea what are the factors mostly creating this phenomenon.”
Get customer buy-in: P8 states that XAI tools are useful in getting customer buy-in and convincing the customers of the model’s decision suggestion: “Will [the customers] be happy with [high accuracy]? Probably. Will they be convinced? Will they be 100% confident that that’s the way they want to do it? Probably not. So it’s all about the buy-in that you get, and SHAP definitely helps.”
Sanity check for model reasoning: One particular case of using XAI for model development is as a sanity check for model reasoning (“If it’s more like an ideation project, then I would say that the explainability analysis [is done] to see […] what is guiding the prediction as well as […] how are my predictions on the test set explained? And […] does it align with the actual value that I have?” - P8).
B.6. Theme area: User perspectives on TDA
The theme area “User perspectives on TDA” groups four themes that emerged in the participant’s answers when discussing TDA.
Usefulness of XAI is use-case dependent: Some participants find XAI only useful in certain use-cases and under certain conditions such as time availability, intuitiveness of the explanation, and the possibility of taking action based on the explanation. When explanations are used for communication with business departments, they do not deliver useful information when the underlying model is not high-performing yet (P3, P8).
XAI is useful: While participants voiced certain conditions for the usefulness of XAI, the overall majority agreed that explanations are useful in their work (P1, P2, P3, P9, P11).
TDA complementary to feature attribution: Participants are overall positive towards TDA explanations. They mentioned that TDA could be complementary to feature attribution explanations and provide context, as feature attribution information is rather local (e.g., comparing feature attribution maps between training and test samples that are highly related to each other).
Different notions of interesting training samples: Within our participants, we find that there are several different notions of what an interesting training sample to a model output could be among the participants. We detail the codes in this theme in Table 2 of the main paper.
@@ B.6. Theme area: User perspectives on TDA
Appendix C Data generation for the survey study
@@
C.1. Use case: Bird classification app
@@
C.2. Use case: Loan approvement recommendation app
@@
C.3. Model error selection
@@
C.4. Relevant training sample selection
@@
Appendix D Survey study themes
The thematic coding analysis of the survey study (N=31) yielded 34 themes corresponding to four theme areas and four unexpected themes (§ 4.5). In the following, we further elaborate on these themes.
D.1. Theme area: Action depends on user’s hypothesis
The theme area “Action depends on user’s hypothesis” groups 21 themes linked to the reasons for participant’s choices in the action axis. We find that the proposed debugging actions of participants are highly dependent on their hypotheses of the model error’s root cause.
[Hypothesis] Mislabeled data: The model makes an error because of mislabeled data in the training set that is highly related to the erroneous test sample.
[Hypothesis] Incomplete training: The model makes an error because it has not finished training, or was trained using non-optimal hyperparameters and training procedures.
[Hypothesis] Spurious correlation: The model makes an error due to a spurious correlation being present in the test image, that the model learnt from the training data.
[Hypothesis] Similar class features: The model makes an error because the erroneous test sample is similar in features to another class in the dataset.
[Hypothesis] Low dataset diversity: The model makes an error because it was not able to generalise from training on a dataset with low diversity in the samples (i.e., samples of a class are all similar, not showing the breadth of the class).
[Hypothesis] Class overlap: The model makes an error because the class of the erroneous test sample is highly overlapping in terms of class-relevant features with another class and distinctive features may not be present in the data.
[Hypothesis] Dataset imbalance: The model makes an error because it was not able to generalise from an imbalanced dataset (i.e., class imbalance).
[Hypothesis] Small dataset size: The model makes an error because it was not able to generalise from a small dataset, instead it overfit it.
[Action] Remove noisy data: To overcome the error, noisy data such as mislabeled data, is removed from the dataset and the model is retrained.
[Action] Changing label to correct: To overcome the error, incorrectly labelled training data is corrected and the model is retrained.
[Action] Fix labelling pipeline for future data: To overcome the error, the labelling and annotation pipeline is inspected to find a root cause for labelling errors.
[Action] Training process engineering: To overcome the error, hyperparameters and training process protocols are adjusted and tuned before retraining.
[Action] Model architecture engineering: To overcome the error, the model architecture is adjusted (e.g. addition of another layer, choosing a different foundational model).
[Action] Feature engineering: To overcome the error, the feature extractors of the data preprocessing pipeline are inspected. This theme was mainly found for the tabular data use case.
[Action] Upweight to focus on specific features: To overcome the error, certain training data samples should be upweighted so that the model learns the features present in these samples better.
[Action] Feature attribution inspection: To overcome the error, the misleading features must be understood first. Feature attribution could be a useful tool.
[Action] Manual inspection: To overcome the error, the training data is manually inspected to find any data artefacts. Participants mentioned that they already do this as part of their model development pipeline.
[Action] Collect specific new data: To overcome the error, new data related to the erroneous training sample should be collected and added to the training dataset. This can improve dataset diversity and help the model generalise better.
[Action] Synthesise/augment data: To overcome the error, new data samples should be synthesised or existing data samples should be augmented further, targeting the parts of training data that are underrepresented. This theme was mentioned often under the assumption that adding completely new data is not possible.
[Action] Resample/reweight data: To overcome the error, the existing training dataset should be resampled, potentially with a specific weighting to overcome class imbalances.
[Action] Collect any new data: To overcome the error, the model needs to learn from more data. Hence, expanding the dataset overall will likely help the model performance.
D.2. Theme area: Individual metric preferences
The theme area “Individual metric preferences” groups four themes and provides insight into why certain metrics are preferred.
[Action/Metric] Label change and probability shows decision boundary: Probabilities are the preferred metric, especially in combination with the label change action, as it gives the user an intuition about the decision boundary between the predicted and ground-truth class.
[Metric] Probabilities are intuitive and show decision boundaries: Probabilities are the preferred metric because they are an intuitive metric bounded between 0-100. Also, they give the user an intuition about the decision boundary between the predicted and ground-truth class.
[Metric] Ground truth probability is the target: The probability of the ground truth class of the test sample is preferred because this is the target metric to optimise when debugging this erroneous test sample.
[Metric] Loss is most informative in the multi-class case: The loss is preferred because it entails information across many classes and is a direct indicator of how well the model is performing overall. This is particularly interesting in multi-class classification cases.
D.3. Theme area: Reliability of explanations
The theme area “Reliability of explanations” groups three themes related to the overarching need for reliability of the attribution scores.
[Action] Removal shows causal effect: The removal action is preferred because represents the causal effect of the inclusion of a specific training sample. This causal effect is assumed to be meaningful and therefore reliable.
[Number of samples] Group size enough for reliable attribution: A group (i.e., more than one training sample) is preferred for training data attribution because considering the effect of a group is more reliable than of single training samples. The group size should be sufficiently large for the attribution score to be meaningful and hence, reliable.
[Number of samples] Small group to prevent reasoning change: While a group of training samples is preferred, the size of the group should be sufficiently small so that the underlying model stays rather consistent given the proposed action (e.g., if the attribution corresponds to the effect of removing samples, this will change the data distribution which affects the model’s learning, too). If the underlying model were to change too much, the attribution scores could be arbitrary and it is not certain whether the model behaviour will stay similar, making the sample-wise debugging process difficult.
D.4. Theme area: Group attribution preferences
The theme area “Group attribution preferences” groups two themes that elaborate on desired sizes of training data groups as explanations. In contrast to the other axes of action and metric, there is a clear preference towards training data group attribution, emphasising a need for methods that serve this task.
[Number of samples] Group for an overview of the error: A group of training samples is preferred in the explanation of the model error to get a better overview of the error, as it is expected that the samples of the group with the highest attribution are related to the model error. One sample is too little for the user to understand the error and propose an action forward based on the explanation.
[Number of samples] Balance between variety and conciseness: While a group of training samples is preferred, the group size should be large enough to show a variety of different training samples related to the model error, while being small enough to be concise and not overwhelm the user.
D.5. Unexpected themes
In addition to the previous theme areas, we found three additional themes which were unexpected. They provide context to the overall analysis, show two predominant mindsets of developers and point to the research direction of grouping erroneous test samples, which has largely been overlooked.
Focus on the model error at hand: The main focus of the user during the debugging process lies in understanding the model error at hand, and identifying a way to fix the error. This is a local focus.
Focus on the overall model performance: The main focus of the user during the debugging process is to optimise the overall model performance, i.e., having the overall model accuracy in mind. This is a global focus.
Grouping erroneous test samples: In the debugging process, users seek to understand the reasons behind the model error. A single test sample is unlikely to be the only instance of erroneous model behaviour, e.g. spurious correlations. This theme describes the need for presenting the user with multiple test instances that the model errs on similarly so that the user will have a better understanding of the error which will enable them to fix it.
Study interface unfamiliar: In the comments, we occasionally found remarks that the participants were unsure about whether they had correctly fulfilled the task, indicating the overall infamiliarity with sample-wise debugging with TDA explanations.

### v2

@@ Acknowledgements.
- Li et al. (2024) Wenjie Li, Jiawei Li, Christian Schroeder de Witt, Ameya Prabhu, and Amartya Sanyal. 2024. Delta-Influence: Unlearning Poisons via Influence Functions. arXiv preprint arXiv:2411.13731 (2024).
- Liao et al. (2020) Q. Vera Liao, Daniel Gruen, and Sarah Miller. 2020. Questioning the AI: Informing Design Practices for Explainable AI User Experiences. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–15. doi:10.1145/3313831.3376590
@@
- Marzi et al. (2024) Giacomo Marzi, Marco Balzano, and Davide Marchiori. 2024. K-Alpha Calculator–Krippendorff’s Alpha Calculator: A user-friendly tool for computing Krippendorff’s Alpha inter-rater reliability coefficient. MethodsX 12 (2024), 102545.
@@
- Nauta et al. (2023) Meike Nauta, Jan Trienes, Shreyasi Pathak, Elisa Nguyen, Michelle Peters, Yasmin Schmitt, Jörg Schlötterer, Maurice van Keulen, and Christin Seifert. 2023. From Anecdotal Evidence to Quantitative Evaluation Methods: A Systematic Review on Evaluating Explainable AI. ACM Comput. Surv. 55, 13s, Article 295 (jul 2023), 42 pages. doi:10.1145/3583558
@@
- Pezeshkpour et al. (2022) Pouya Pezeshkpour, Sarthak Jain, Sameer Singh, and Byron Wallace. 2022. Combining Feature and Instance Attribution to Detect Artifacts. In Findings of the Association for Computational Linguistics: ACL 2022. Association for Computational Linguistics, Dublin, Ireland, 1934–1946. doi:10.18653/v1/2022.findings-acl.153
- Piorkowski et al. (2023) David Piorkowski, Inge Vejsbjerg, Owen Cornec, Elizabeth M. Daly, and Öznur Alkan. 2023. AIMEE: An Exploratory Study of How Rules Support AI Developers to Explain and Edit Models. Proc. ACM Hum.-Comput. Interact. 7, CSCW2, Article 255 (Oct. 2023), 25 pages. doi:10.1145/3610046
@@
- Rong et al. (2023) Yao Rong, Tobias Leemann, Thai-Trang Nguyen, Lisa Fiedler, Peizhu Qian, Vaibhav Unhelkar, Tina Seidel, Gjergji Kasneci, and Enkelejda Kasneci. 2023. Towards Human-Centered Explainable AI: A Survey of User Studies for Model Explanations. IEEE Trans. Pattern Anal. Mach. Intell. 46, 4 (nov 2023), 2104–2122. doi:10.1109/TPAMI.2023.3331846
@@
- Sambasivan et al. (2021) Nithya Sambasivan, Shivani Kapania, Hannah Highfill, Diana Akrong, Praveen Paritosh, and Lora M Aroyo. 2021. “Everyone wants to do the model work, not the data work”: Data Cascades in High-Stakes AI. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (¡conf-loc¿, ¡city¿Yokohama¡/city¿, ¡country¿Japan¡/country¿, ¡/conf-loc¿) (CHI ’21). Association for Computing Machinery, New York, NY, USA, Article 39, 15 pages. doi:10.1145/3411764.3445518
- Schioppa et al. (2022) Andrea Schioppa, Polina Zablotskaia, David Vilar, and Artem Sokolov. 2022. Scaling Up Influence Functions. Proceedings of the AAAI Conference on Artificial Intelligence 36, 8 (Jun. 2022), 8179–8186. doi:10.1609/aaai.v36i8.20791
- Shen et al. (2023) Hua Shen, Chieh-Yang Huang, Tongshuang Wu, and Ting-Hao Kenneth Huang. 2023. ConvXAI: Delivering Heterogeneous AI Explanations via Conversations to Support Human-AI Scientific Writing. In Companion Publication of the 2023 Conference on Computer Supported Cooperative Work and Social Computing (Minneapolis, MN, USA) (CSCW ’23 Companion). Association for Computing Machinery, New York, NY, USA, 384–387. doi:10.1145/3584931.3607492
@@
- Wang et al. (2025a) Andrew Wang, Elisa Nguyen, Runshi Yang, Juhan Bae, Sheila A McIlraith, and Roger Grosse. 2025a. Better Training Data Attribution via Better Inverse Hessian-Vector Products. arXiv preprint arXiv:2507.14740 (2025).
@@
- Wang et al. (2024) Fulton Wang, Julius Adebayo, Sarah Tan, Diego Garcia-Olano, and Narine Kokhlikyan. 2024. Error discovery by clustering influence embeddings. In Proceedings of the 37th International Conference on Neural Information Processing Systems (New Orleans, LA, USA) (NIPS ’23). Curran Associates Inc., Red Hook, NY, USA, Article 1809, 13 pages.
- Wang et al. (2025b) Jiachen T Wang, Dawn Song, James Zou, Prateek Mittal, and Ruoxi Jia. 2025b. Capturing the Temporal Dependence of Training Data Influence. In The Thirteenth International Conference on Learning Representations.
@@
- Williams (2021) Oyindamola Williams. 2021. Towards human-centred explainable ai: A systematic literature review. Master’s Thesis (2021).
- Wolf (2019) Christine T Wolf. 2019. Explainability scenarios: towards scenario-based XAI design. In Proceedings of the 24th International Conference on Intelligent User Interfaces. 252–257.
@@ Appendix A Inclusion criteria: High-risk application areas
Appendix B Data generation for the scenario-based interactive user study
@@
B.1. Use case: Bird classification app
@@
B.2. Use case: Loan approvement recommendation app
@@
B.3. Model error selection
@@
B.4. Relevant training sample selection
@@
Appendix C Extended background: Training Data Attribution Methods
We provide a high-level overview of TDA methods cited in Table 4 of the main paper and how they relate to the attribution types. Specifically, we elaborate on influence functions methods (§ C.1), methods that trace the training process (§ C.2), methods that unroll the training process (§ C.3) and group influence methods (§ C.4).
C.1. Influence Function Methods
Influence functions (IF), first introduced in robust statistics (Hampel, 1974), are one of the main types of methods in TDA. Instead of costly retraining without a training sample to quantify ’s influence on the model’s behaviour via Equation 1, (IF) approximate this quantity by:
where uses a Newton-like step to estimate the effect of a small perturbation on the model’s parameters, where the perturbation is the exclusion of . For a full derivation, we refer to Koh and Liang (2017) or Bae et al. (2022). Intuitively, influence functions approximate the parameter landscape around the parameters , and compute the update step taken in the direction of removing , to estimate its effect on the loss at .
IF require approximations to be applicable to deep models.
Since equation 3 requires the inversion of the parameter Hessian (), it also requires the objective function (i.e., the loss) to be twice-differentiable and the Hessian to be positive definite, to ensure that exists. However, these conditions generally do not hold for deep models. Moreover, the Hessian is a large and costly matrix to invert, so IF algorithms usually do not explicitly estimate but implicitly through a matrix-vector product. They estimate the damped inverse Hessian Vector Product (iHVP) , where is the damping parameter and is the identity matrix. Damping improves the numerical stability of the iHVP estimation. In the following, we use the terms damped iHVP and iHVP interchangeably.
IF methods generally differ in iHVP approximation.
Different TDA methods propose different approximations of the iHVP: Koh and Liang (2017), who first presented IFs for deep models, estimate the iHVP using an iterative estimator called Linear time Stochastic Second-Order Algorithm (LiSSA) (Agarwal et al., 2017). Guo et al. (2021) also use LiSSA in their algorithm, but propose an additional speed up for the task of identifying the most influential training points for a test point by first retrieving candidate with a k-nearest neighbour search in the model’s embedding space. Wang et al. (2025a) improve the iHVP and therefore influence estimation by using the Eigenvalue-corrected Kronecker-Factored Approximate Curvature (EKFAC) (Grosse et al., 2023) as a preconditioning matrix on the iterative estimator. Alternatively, Schioppa et al. (2022) propose to speed up the approximation of the iHVP by projecting the computation into a lower-dimensional space spanned by the dominant eigenvectors of with Arnoldi iterations (Arnoldi, 1951). Also Park et al. (2023) and Choe et al. (2024) leverage a projection into lower dimensions to speed up the iHVP estimation, where the former uses random projections while the latter makes use of gradient structures. Grosse et al. (2023) propose to estimate the utilising Eigenvalue-corrected Kronecker Factored Approximate Curvature (EKFAC), allowing for a faster computation of the influence function since it is not an iterative method.
Influence function research is active and has been gaining traction in recent years. The main focus lies on finding an accurate estimation of the loss difference on a single test sample when a single training sample is removed.
C.2. Methods That Trace The Training Process
In 2020, Pruthi et al. (2020) presented a different way to compute training data attribution that traces a training sample’s contribution to the model performance on a test sample throughout training. The objective definition of TDA is slightly different from the change after leave-one-out retraining (Eq. 1):
TDA is hence defined as the sum of the loss difference on at each training step with across the full training.
In practice, computing this idealised notion of TDA is also prohibitive across all training samples, so Pruthi et al. (2020) propose to approximate equation 4 with a sum of gradient dot products across model checkpoints recorded during training:
where is the learning rate at checkpoint . We note that in principle, each dot product corresponds to the formulation of influence functions (Eq. 3) where the inverse Hessian is approximated by the identity matrix.
TracIn (Pruthi et al., 2020) offers an easy-to-compute TDA method and objective. It is focused on the relationship between a single training and test sample, measured on the loss changes across training, whenever the training sample is seen during training.
C.3. Methods That Unroll The Training Process
In a similar yet different spirit to tracing the training process as in TracIn (Pruthi et al., 2020), another line of approaches to computing TDA are methods that are based on unrolling the training process (Hara et al., 2019; Bae et al., 2024; Wang et al., 2025b; Ilyas and Engstrom, 2025), sometimes also referred to as dynamic TDA methods (Hammoudeh and Lowd, 2024). Unrolling the training process results in a computational graph which can be traced backwards to compute attribution. Hence, these methods also consider the training procedure, but differentiate through it to approximate the change in model parameters with leave-one-out retraining (cf. equation 1) (Hara et al., 2019).
Since differentiating through each training step is computationally expensive, methods often group several steps in training segments and base their approximation on a sum of influence function-like computations per segment. Similar to influence functions, Hara et al. (2019) leverage the implicit function theorem to compute the Hessian vector product, and Bae et al. (2024) make use of EKFAC approximations of the Hessian (Grosse et al., 2023) for the influence-like computation. Ilyas and Engstrom (2025) propose a different approach using metagradients (Engstrom et al., 2025) to explicitly compute the influence function for single-model training data attribution, fixing sources of training process stochasticity such as batch ordering and model initialisation. While this approach achieves near-optimal performance, it is very costly.
As unrolling-based methods, like influence functions, aim at capturing the effect of leave-one-out retraining, they provide methods to consider deep model training specificities (e.g., learning rate schedules) and generally target the attribution of model behaviour measured in loss on a single test sample to a single training sample when it is removed from training.
C.4. Group Attribution Methods
TDA methods generally address the individual attribution of the model behaviour to a training sample . Another line of work studies methods for quantifying the group attribution of to a set of training samples . The TDA objective is analogous to individual attribution, and is generally quantified by the loss difference after leave-some-out retraining:
Like individual attribution, computing equation 6 directly is computationally prohibitive when the objective is to identify the group of training samples with the largest attribution. Hence, group TDA work studies methods to approximate this quantity efficiently.
Group attribution assuming linearity
An approach that limits the computational cost makes the core assumption that group attribution is linear: In other words, the attribution score of a group of training samples is defined as the sum of the individual attribution scores of the training samples . Koh et al. (2019) found that interaction effects between training samples of a group can be neglected, and that the linearity assumption is sufficiently accurate for group TDA estimation. As a result, all individual TDA methods, like influence functions, can also be used to estimate group TDA. Moreover, the common evaluation metric linear datamodeling score (Park et al., 2023) used to evaluate the performance of TDA methods is based on the assumption of linearity.
Group attribution considering interaction effects
Group TDA methods that assume linearity have the same computational cost as individual TDA methods, but ignore any potential interaction effects between training samples of a group. While Koh et al. (2019) found the interaction effects to be minimal and therefore negligible, a different work by Basu et al. (2020) improves the accuracy of influence functions in capturing equation 6 by extending influence functions with a second-order term to capture interaction effects. Shapley value-based approaches like Data Shapley (Ghorbani and Zou, 2019; Lin et al., 2024) also consider interaction effects by default. Hu et al. (2024) propose an adaptive algorithm that iterates over the training dataset and gradually identifies the most influential subset for a given model behaviour.